Circular No. 7101

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

RX J0052.1–7319 AND XTE J0111.2–7317 G. L. Israel and L. Stella, Osservatorio Astronomico, Rome; S. Covino

and S. Campana, Osservatorio Astronomico di Brera, Milan; and S. Mereghetti, Istituto Fisica Cosmica e Tecnologie Relative, CNR, Milan, report the discovery with the Danish 1.54-m telescope at the European Southern Observatory of the optical counterparts of the two recently discovered x-ray pulsar transients in the Small Magellanic Cloud: “We have identified the likely optical counterpart of RX J0052.1–7319 (IAUC 7081): a B-type star with R = 14.54 ± 0.03 and V − R = +0.08 ± 0.04 that is located at α = 0^{h}52^{m}14^{s}.0, δ = −73^{o}19′18′′ (equinox 2000.0; estimated uncertainty 1′′). A 1000-s optical spectrum taken on Jan. 19 (range 350–900 nm, resolution 1.1 nm) shows an Hα emission line (EW = 1.2 nm), clearly indicating the Be nature of the star. Within the 10′′-radius x-ray error circle, there is another object (R = 16.05 ± 0.05) that does not show any emission lines. The field of XTE J0111.2–7317 (IAUC 7048), a currently active x-ray pulsar transient, was observed with the same instrument on Jan. 19, revealing two bright stars within the 30′′ error circle (IAUC 7062); the fainter one is a B-type star (R = 15.29 ± 0.03, V − R = +0.06 ± 0.04) located at α = 1^{h}11^{m}08^{s}.4, δ = −73^{o}16′46′′ (1′′ uncertainty). A 1000-s spectrum (range 350–900 nm, resolution 1.4 nm) revealed strong Hα and Hβ emission lines (EW = 2.1 and 0.1 nm, respectively), indicating that this source is the likely counterpart of the x-ray transient. The brighter star (R = 14.30 ± 0.03) does not show any emission lines. No other objects brighter than R = 17.3 are found within the error circle. Another B-type star (R = 14.55 ± 0.03, V − R = +0.05 ± 0.04), located at α = 1^{h}11^{m}25^{s}.9, δ = −73^{o}17′27′′ (30′′ outside the x-ray error circle), shows strong Hα and Hβ emission lines (EW = 3.6 and 0.3 nm, respectively)”.

SUPERNOVA 1998co IN NGC 7131 Corrigendum. The following position by E. Prosperi from a CCD

frame taken on Jan. 26.736 UT corrects that on IAUC 7083, line 4: α = 21^{h}47^{m}36^{s}.24, δ = −13^{o}10′52′′.3.

SUPERNOVA 1999B IN UGC 7189 Unfiltered CCD magnitude estimates: Jan. 21.716 UT, 17.3 ± 0.2

(L. Kiss, K. Sárneczky, and E. Barát, Piszkéstető, Konkoly Observatory); 21.88, 17.3 (S. Moretti, S. Tomaselli, and G. Cortini, Forlí, Italy).

1999 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green